How much heat is lost when a 56g glass of 80 °C water cools down to 0 °C.
1
Expert's answer
2020-04-20T15:05:44-0400

Solution:

To calculate the amount of heat entering or leaving a system, the equation Q = mcΔT is used.

m = mass (in grams)

c = specific heat capacity (J/g°C)

ΔT = change in temperature (°C).


Specific heat of water is 4.18 J/goC

ΔT = (Tfinal - Tinitial) = 0°C - 80°C = -80°C

Then,

Q = mcΔT = (56 g) * (4.18 J/g°C) * (-80°C) = -18726.4 J = -18.73 kJ.

Q = -18.73 kJ.

The negative Joule value means heat is being lost in an exothermic reaction.


Answer: Q = -18.73 kJ.